the sum of three consecutive numbers = 258 What are these numbers
Answer by checkley77(12844)

Let x, x+1 & x+2 be the 3 numbers.
x+x+1+x+2=258
3x=258-3
3x=255
x=255/3
x=85 ans. for the smaller number.
85+1=86 ans. for the middle number.
85+2=87 ans. for the larger umber.
Proof:
85+86+87=258
258=258